3. Given the following, find Q. P = 200 - 3Q MC = 3Q FC = 0
Answer #1

For a competitive firm P=MC

200-3Q=3Q

6Q=300

Q=50

For a monopoly/monopolistic firm MR=MC

TR=PQ

TR=200Q-3Q^2

MR=d(TR)/dQ

MR=200-6Q

At equilibrium:

200-6Q=3Q

9Q=200

Q=22.23

Q depends on the type of market.